# Tau Tridents at Accelerator Neutrino Facilities

### v1

@@ V Conclusions
The neutrino trident process provides an important precision test of the SM electroweak interactions as well as a powerful probe for BSM physics. The next-generation accelerator neutrino experiment DUNE is in an excellent position to observe these rare SM processes for the first time. Among all trident channels, those involving tau leptons are particularly interesting, because any potential appearance of tau leptons at DUNE ND has been associated with BSM physics in the literature. Therefore, it is of paramount importance to delineate all potential SM backgrounds for tau events.
In this paper, we have identified one such SM source of tau lepton production at DUNE ND, namely, tau tridents. To this effect, we performed for the first time a detailed study of tau tridents at DUNE ND, and contrary to the common lore, found that they could yield a non-negligible event rate, especially with the tau-optimized flux configuration. We calculated the tau trident cross section using the full scattering in order to obtain an accurate prediction of the event rate – significantly higher than the cross section calculated with the equivalent photon approximation. We found an intricate interplay between coherent and incoherent scattering which is distinct for tau tridents from electron and muon tridents. Lastly, we have identified a promising venue to enhance the poor signal-to-background ratio using kinematic variables, potentially leading to the observability of tau tridents for the first time at the DUNE experiment. We hope that these findings will motivate a dedicated experimental feasibility study of tau tridents at accelerator neutrino facilities.
@@ Acknowledgements.
We would like to thank Wolfgang Altmannshofer and Bei Zhou for useful discussions. We also thank Toni Mäkelä and Sebastian Trojanowski for useful correspondence on the FASER trident analysis. This manuscript has been authored in part by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work of BD was partly supported by the U.S. Department of Energy under grant No. DE-SC 0017987. DLG is supported by an MCSS Graduate Fellowship.Appendix A Form Factors
Here we give the expressions for the dimensionless flux factors appearing in Eq. (II). In the coherent regime, these have the form [18]
where is the incoming neutrino energy and is the nuclear mass. is the nuclear form factor. For argon and iron we have used
@@
In the incoherent (diffractive) regime, the flux factors appearing in Eq. (II) have the form [18]
where is the proton or neutron mass, and are the nucleon form factors, defined as [18]
where . The functions and here can be related to the electric and magnetic form-factors of the nucleons as
@@
Appendix B Neutrino Fluxes
@@
Appendix C Trident Events at Other Accelerator/Collider Neutrino Experiments
@@
Appendix D Tau Trident Event Distributions at DUNE

### v2

@@ V Conclusions
Any potential appearance of tau leptons at near detectors in accelerator-neutrino experiments has been associated with BSM physics in the literature. Therefore, it is of paramount importance to delineate all potential SM backgrounds for tau events. In this work, we identified one such SM source of tau lepton production at DUNE-ND, namely, tau tridents. To this effect, we performed for the first time a detailed study of tau tridents at DUNE-ND, and contrary to the common lore, found that they could yield a non-negligible event rate, especially with the tau-optimized flux configuration. We calculated the tau trident cross section using the full scattering in order to obtain an accurate prediction of the event rate and found an intricate interplay between coherent and incoherent scattering which is distinct for tau tridents from electron and muon tridents. Lastly, we have identified a promising venue to enhance the poor signal-to-background ratio using kinematic variables, potentially leading to the observability of tau tridents for the first time in the DUNE experiment. We hope that these findings will motivate a dedicated experimental feasibility study of tau tridents at accelerator neutrino facilities.
@@ Acknowledgements.
We would like to thank Wolfgang Altmannshofer, Alex Sousa and Bei Zhou for useful discussions. We also thank Toni Mäkelä and Sebastian Trojanowski for useful correspondence on the FASER trident analysis. This manuscript has been authored in part by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work of BD was partly supported by the U.S. Department of Energy under grant No. DE-SC 0017987. DLG is supported by an MCSS Graduate Fellowship.Appendix A Equivalent Photon Approximation Comparison
Here we remark on an alternative approach for calculating the component of the cross section in Eq. (2), utilizing the equivalent photon approximation (EPA) [52]. To date, the only tau trident estimates for DUNE are given in Ref. [18], which employed the EPA, and deemed these processes inaccessible at DUNE energies.222Note also that the projected neutrino fluxes used in Ref. [18] have since been updated to those we employ in this work, and so the exact projected event numbers should not be directly compared. In the EPA, the infrared pole of the propagator for on-shell photon exchange between the lepton and nucleus is assumed to yield a dominant contribution to the cross section. As , the transverse cross section approaches that of an on-shell photon, and the longitudinal component vanishes. Thus, when applying the EPA, the longitudinal contributions in Eq. (2) are omitted, and the total cross section scales roughly as [52, 18], where is the Fermi constant.
The EPA method gives a good estimate for some neutrino trident cross sections when coherent scattering dominates [53]. It overestimates the cross sections of the dielectron trident, but is a good approximation for the coherent contribution to dimuon tridents [19, 26]. Ref. [19] establishes that for accelerator neutrino energies, EPA is not valid for tridents in regions where the virtual photon contributions are considerable, and where the incoherent scattering contributions are non-negligible – as we see is true for tau-containing tridents from our results in the main text.
Appendix B Form Factors
Here we outline the utilized nuclear and nucleon form factors, used for the calculation of coherent and diffractive neutrino trident scattering, respectively. For nuclear form factors, we highlight a series of benchmark target nuclei relevant not only for DUNE but also for the other experiments for which we provide event estimates in Appendix D.
B.1 Nuclear form factors
For argon and iron, we have used
@@ B.1 Nuclear form factors
B.2 Nucleon form factors
@@
Appendix C Neutrino Fluxes
@@
Appendix D Trident Events at Other Accelerator/Collider Neutrino Experiments
@@
Appendix E Tau Trident Event Distributions at DUNE